OpenStudy (anonymous):

Use the definition or identities to find the exact value of the six trigonometric functions 4) tan θ = 3

OpenStudy (anonymous):

tan(@) = sin(@)/cos(@) ? or work with c^2 = a^2 + b^2 tan(@) = a/c etc.